# Always skip attention

### v2

@@ Abstract
We highlight a curious empirical result within modern Vision Transformers (ViTs). Specifically, self-attention catastrophically fails to train unless it is used in conjunction with a skip connection. This is in contrast to other elements of a ViT that continue to exhibit good performance (albeit suboptimal) when skip connections are removed. Further, we show that this critical dependence on skip connections is a relatively new phenomenon, with previous deep architectures (e.g., CNNs) exhibiting good performance in their absence. In this paper, we theoretically characterize that the self-attention mechanism is fundamentally ill-conditioned and is, therefore, uniquely dependent on skip connections for regularization. Additionally, we propose Token Graying(TG)—a simple yet effective complement (to skip connections) that further improves the condition  [^1] We remind readers that the condition number of a full-rank matrix is defined by the ratio of the largest singular value to the smallest singular value. A lower condition number implies better condition. of input tokens. We validate our approach in both supervised and self-supervised training methods.
Figure 1: We demonstrate a curious phenomenon: removing skip connections from the Self-Attention Block (SAB) versus the Feedforward Network (FFN) in Vision Transformer (ViT) models results in different performance drop. (a) Model architecture of ViT consisting of a SAB, and a FFN in each layer. ① represents the output of the self-attention transformation, while ③ represents the output of the MLP transformation. ② and ④ represent identity mappings, referred to as skip connections, in the SAB and FFN, respectively. (b) This figure demonstrates the condition number of the different output embeddings for each hidden layer. Note this model of ViT-Tiny is trained in a typical manner with skip connections on both SAB and FFN blocks (i.e., ① + ② + ③ + ④). It shows that the SAB output embedding without skip connection is poorly conditioned (①), with a condition number significantly higher than that of the other three configurations involving ViT blocks. (c) Classification results on three different datasets using the ViT-Tiny model under three different architectural configurations at train time. The result shows the effects of removing skip connections at train time on the SAB and FFN blocks (either ② or ④) in all layers of ViT-Tiny. We observe that removing ② results in a catastrophic performance drop, whereas removing ④ (as predicted) has a modest impact on performance. This difference becomes more pronounced as the scale of the dataset increases.
@@ 1 Introduction
Vision Transformers (ViTs) have demonstrated impressive performance on various computer vision tasks, such as object detection, semantic segmentation, video understanding, visual-language learning, and many others [14, 27, 12, 9]. Much of this success is credited to the self-attention mechanism—or Self-Attention Block (SAB)—which enables ViTs to selectively focus on relevant parts of the input sequence while generating each element of the output sequence. Another critical element is the MLP—or Feedforward Networks (FFN)—which facilitates intra-token communication across channel dimensions (see Fig. 1). Finally, an identity mapping—commonly called a skip connection—is included in both the SAB and FFN to ensure good performance.
We argue that the Jacobian of the SAB is disproportionately ill-conditioned compared to other components, notably the FFN block. Poor Jacobian condition is fundamentally detrimental to gradient descent training [15], making convergence and stability challenging. An obvious strategy in this regard is to directly model the Jacobian of the entire network. This is problematic in practice due to sheer size of modern transformer networks. So we make the following simplifying assumptions:
@@
Although we do not formally prove these assumptions, we indirectly validate them through our extensive empirical analysis. First, we show that better conditioning of a sub-block’s output embeddings leads to improved empirical performance across several ViT benchmarks. Second, we empirically demonstrate that the Jacobian condition of the sub-block is improved( Appendix C).
Skip in transformers
@@
Skip connections have become an indispensable component in modern transformer neural networks [23] and have been empirically proven to be a de facto component. In Fig. 1 (b), we take one pre-trained regular ViT-Tiny model and measure the condition of the SAB and FFN outputs, both with and without skip connections. We observe that the SAB output embedding without skip connections is highly ill-conditioned, with a condition number around $e^{6}$. In contrast, the other three configurations have relatively low condition numbers, around $e^{3}$. In other words, the output embedding of SAB without skip connections has significantly worse condition than the other configurations. Consequently, we train the models from scratch with three different configurations: the standard SAB and FFN with skip connections, SAB without skip connections, and FFN without skip connections.
@@
In Fig. 1 (c) we observe—for a ViT-Tiny model trained upon the CIFAR-10 dataset—that classification accuracy drops modestly ($2\%$) when skip connections are removed in the FFN, but retained in the SAB. Surprisingly, however, the model performance becomes near catastrophic—with a $22\%$ drop in accuracy—when skip connections are removed from the SAB while retained in the FFN. Further, as dataset size increases, models without skip connections in the SAB tend to degrade significantly in performance and at a larger scale (e.g., Tiny-ImageNet) are not trainable at all. This empirical result, to our knowledge, has not been reported previously in the literature. Additionally, the training loss of three configurations is shown in Fig. 3. The results demonstrate that the SAB without skip connections—which is significantly ill-conditioned—converges much more slowly than the other two configurations and diverges after 30 epochs. Interestingly, this critical dependence on skip connections is unique to transformers.
Skip in ConvMixer.
@@
To further highlight the poor conditioning of the SAB, we conducted a similar experiment on a modern CNN ConvMixer [22]. ConvMixer has a similar architecture and competitive performance to ViT with the exception that self-attention is replaced by a convolution block to spatially mix tokens. Unlike ViT, training ConvMixer without skip connections still achieves competitive performance, as shown in Fig. 2. This contrast further suggests that the SAB introduces an extreme conditioning issue, making skip connections indispensable for stable training in ViTs.
Contributions.
@@
    We present a proposition that characterizes why SAB output embedding without skip connection is fundamentally ill-conditioned, which challenges training convergence and stability (Sec. 4.1).
@@
    A theoretical analysis on the role of skip connection within the SAB is undertaken. We demonstrate that it significantly improves the condition of the block’s output embedding, enhancing stability and performance (Sec. 4.2).
@@
    Finally, we propose a novel approach—Token Greying (TG)—to better pre-condition the input tokens. This step, when used in conjunction with conventional skip connections—improves ViT in both supervised and self-supervised settings (Sec. 5).
@@ 2 Related work
Initially proposed by Vaswani et al. [23], self-attention has become central to transformer architectures because it captures long-range dependencies by dynamically weighting interactions between tokens. Building on this foundation, transformer-based large language models—such as Llama [5] and Deepseek [10]—have significantly advanced language understanding. In computer vision, researchers introduced Vision Transformers (ViTs) [4] for image classification, achieving superior results compared to traditional convolutional neural networks. Moreover, researchers have trained Diffusion Transformers [16], replacing the commonly used U-Net backbone with a transformer that operates on latent patches and inherits the excellent scaling properties of the transformer model class. Similarly, self-attention mechanisms have been successfully applied in speech recognition and physics-informed neural networks, further demonstrating their versatility across diverse applications [2, 26].
@@
He et al. [8] introduced skip connections in ResNet to make deep networks easier to optimize and to enhance accuracy by enabling significantly increased network depth. Since their introduction, skip connections have been extensively studied and applied across various Convolutional Neural Network (CNN) architectures. Building on these ideas, Veit et al. [24] offer a novel interpretation of residual networks, demonstrating that they can be viewed as a collection of multiple paths of varying lengths, which contributes to the network’s flexibility and resilience. Furthermore, Hayou et al. [6] addresses gradient stability in deep ResNets, noting that while skip connections alleviate the issue of vanishing gradients, they may also introduce gradient explosion at extreme depths. To counter this, they propose scaling the skip connections according to the layer index, thereby stabilizing gradients and ensuring network expressivity even in the limit of infinite depth. Recently, in Table 3 of [11], the authors reported that VGG—a model without skip connections—achieves comparable results to ResNet—a model with skip connections—on modern datasets. To the best of our knowledge, only one paper challenges skip connection in transformer architectures. In [7], the authors attempted to train deep transformer networks without using skip connections, achieving performance comparable to standard models. However, this approach requires five times more iterations, and the reasons why skip connections are crucial for self-attention-based transformers remain unresolved.
@@ 3 Preliminary
In this section, we outline the definitions of the self-attention blocks used in the Vision Transformer architecture and establish the mathematical notation for the quantities that will be utilized in the following sections. For more detailed information on transformers, readers are encouraged to consult [4].
A Vision Transformer architecture consists of L stacked self-attention blocks (SABs) and feedforward networks (FFNs). An identity mapping—commonly referred to as a skip connection—is applied to connect inputs and outputs of transformation in both SAB and FFN.
@@
where self-attention output embedding is
@@
where $\mathbf{Q}=\mathbf{X}_{\text{in}}\mathbf{W}_{\text{Q}}$, $\mathbf{K}=\mathbf{X}_{\text{in}}\mathbf{W}_{\text{K}}$ and $\mathbf{V}=\mathbf{X}_{\text{in}}\mathbf{W}_{\text{V}}$ with learnable parameters $\mathbf{W}^{\text{Q}},\mathbf{W}^{\text{K}},\text{and }\mathbf{W}^{\text{V}}\in\mathbb{R}^{d\times d}$. $\eta(\cdot)$ is an activation function and is set by default as $\operatorname*{softmax}$ function in [4]. Additionally, some studies [19, 17, 18] have explored the possibilities of using alternative activation functions. In particular, good performance is achieved by using linear attention with an appropriate scale. It allows us to characterize why the SAB output embedding without skip connection is ill-conditioned.
@@
where $\|\cdot\|_{2}$ is the matrix operator norm induced by the Euclidean norm. We use the Euclidean norm in this paper. $\mathbf{A^{\dagger}}$ is the pseudo inverse of the matrix $\mathbf{A}$. $\sigma_{\text{max}}$ and $\sigma_{\text{min}}$ are the maximal and minimal singular values of $\mathbf{A}$ respectively. We will use Eq. 4 as a metric for analysis in future sections.
@@ 4.1 Self-attention is ill-conditioned
Assume $\mathbf{X}\in\mathbb{R}^{n\times d}$,$\mathbf{W_{Q}}$, and $\mathbf{W_{K}}$ and $\mathbf{W_{V}}\in\mathbb{R}^{d\times d}$ have entries that are independently drawn from a distribution with zero mean. The condition number of the SAB output embedding (without skip-connection) can be expressed as:
@@
Fig. 4 illustrates our theoretical analysis by showing that the linear self-attention operation increases the condition number of embeddings from $e^{3}$ to $e^{6}$, and we can empirically extend this analysis to softmax self-attention. The output embedding of the softmax SAB without skip connections is also significantly ill-conditioned.
@@ 4.2 Skip connection improves condition
Previous works [7, 13, 3] have demonstrated that transformers, without skip connections, converge very slowly or may even become un-trainable. In [7], the authors show that without skip connections, output embeddings of the SAB quickly converge to rank 1. They hypothesize that this is due to the inability of the transformer to propagate signals through the self-attention mechanism. This rank-1 output embedding implies an infinite condition number. The authors introduce an inductive bias into the self-attention mechanism to enable what they refer to as “better signal propagation”. The approach matched the performance of the vanilla models, but at the cost of being five times slower. We speculate that the success of the method could be directly attributed to its ability to improve the condition of the self-attention output embeddings.
This inspires us to speculate that the true role of skip connection may be to regularize the condition of the self-attention output embeddings
@@
Proposition 4.2 shows that during training, the condition number of the SAB output embedding has a much lower bound than that of the SAB output embedding without skip connections. Furthermore, with other conditions being equal, better condition of SAB can contribute to faster convergence and more stable gradient updates, reducing the risk of gradient explosion or vanishing. This ensures the signal can propagate well during the training process in the Vision Transformers.
@@
As shown in Fig. 4, empirical results demonstrate that in both the linear and $\operatorname*{softmax}$ self-attention cases, adding skip connections significantly improves the condition of output embeddings.
@@ 5 Methodology
1$\triangleright$ Input: training data, amplification coefficient $\epsilon\in(0,1]$
2$\triangleright$ Training:
@@
5:     $\mathbf{U\Sigma V^{T}\leftarrow X}$triangleright Compute SVD
6:     $\mathbf{\tilde{\Sigma}\leftarrow\frac{\Sigma}{\text{max}(\Sigma)}}$triangleright Normalize elements $\in(0,1]$
7:     $\mathbf{\tilde{\Sigma}\leftarrow\tilde{\Sigma}^{\epsilon}}$triangleright Amplify
8:     $\mathbf{\tilde{X}\leftarrow U\tilde{\Sigma}V^{T}}$triangleright SVD reconstruct
9:     Patch Embedding: $\mathbf{Z}\leftarrow\text{PatchEmbed}(\tilde{\mathbf{X}})$

### v3

@@ Abstract
We highlight a curious empirical result within modern Vision Transformers (ViTs). Specifically, self-attention catastrophically fails to train unless it is used in conjunction with a skip connection. This is in contrast to other elements of a ViT that continue to exhibit good performance (albeit suboptimal) when skip connections are removed. Further, we show that this critical dependence on skip connections is a relatively new phenomenon, with previous deep architectures (e.g., CNNs) exhibiting good performance in their absence. In this paper, we theoretically characterize that the self-attention mechanism is fundamentally ill-conditioned and is, therefore, uniquely dependent on skip connections for regularization. Additionally, we propose Token Graying (TG), a simple yet effective complement (to skip connections) that further improves the condition of input tokens. We validate our approach in both supervised and self-supervised training methods.
Figure 1: Removing skip connections from the Self-Attention Block (SAB) versus the Feedforward Network (FFN) in Vision Transformer (ViT) models results in different performance drop. (a) Model architecture of ViT consisting of a SAB, and a FFN in each layer. ① represents the output of the self-attention transformation, while ③ represents the output of the MLP transformation. ② and ④ represent identity mappings, referred to as skip connections, in the SAB and FFN, respectively. (b) The condition number of the different output embeddings for each hidden layer. A ViT-Tiny model is trained in a typical manner with skip connections on both SAB and FFN blocks (i.e., ① + ② + ③ + ④). It shows that the SAB output embedding without skip connection is poorly conditioned (①), with a condition number significantly higher than that of the other three configurations involving ViT blocks. (c) Classification results on three different datasets using the ViT-Tiny model under three different architectures at train time. The result shows the effects of removing skip connections at train time on the SAB and FFN blocks (either ② or ④) in all layers of ViT-Tiny. We observe that removing ② results in a catastrophic performance drop, whereas removing ④ (as predicted) has a modest impact on performance. This difference becomes more pronounced as the scale of the dataset increases.
@@ 1 Introduction
Vision Transformers (ViTs) have demonstrated impressive performance on various computer vision and robotic tasks, such as object detection, semantic segmentation, video understanding, visual-language learning, and many others [19, 33, 17, 12, 8]. Much of this success is attributed to the self-attention mechanism, commonly referred to as Self-Attention Block (SAB), which allows ViTs to selectively attend to relevant parts of the input sequence when generating each element of the output sequence. Another essential component is the MLP, or Feedforward Networks (FFN), which facilitates intra-token communication across channel dimensions (see Fig.˜1). Finally, identity mappings, commonly referred to as skip connections, are incorporated into both the SAB and FFN to enhance model performance.
We argue that the Jacobian of the SAB is disproportionately ill-conditioned compared to other components, notably the FFN block. A poorly conditioned  [^1] The condition number of a full-rank matrix is defined by the ratio of the largest singular value to the smallest singular value. A lower condition number implies better condition. Jacobian is fundamentally detrimental to gradient descent training  [20, 25, 24, 13, 1], impeding both convergence and stability. A natural strategy to address this issue is to explicitly model the Jacobian of the entire network. However, this approach is impractical due to the sheer scale of modern transformer architectures. To proceed, we introduce the following simplifying assumptions:
@@
Although we do not provide formal proofs for these assumptions, we offer indirect validation through extensive empirical analysis. First, we show that better conditioning of a sub-block’s output embeddings consistently correlates with improved empirical performance across multiple ViT benchmarks. Second, we demonstrate that this improvement in embedding condition is accompanied by a measurable enhancement in the sub-block’s Jacobian condition
@@
Skip in Transformers. Skip connections have become an indispensable component in modern transformer architectures [29] and have been empirically proven to be a de facto component. In Fig. 1 (b), we analyze a pre-trained ViT-Tiny model by measuring the condition numbers of the SAB and FFN output embeddings, both with and without skip connections. Our observations reveal that the SAB output embedding becomes highly ill-conditioned in the absence of skip connections, with a condition number approaching $e^{6}$. In contrast, the other three configurations have relatively low condition numbers, around $e^{3}$. In other words, the SAB output embedding without skip connections has significantly worse condition than the other configurations. To further investigate the impact of skip connections, we train the ViT-Tiny models from scratch under three configurations: the standard SAB and FFN with skip connections, SAB without skip connections, and FFN without skip connections.
@@
In Fig. 1 (c), we observe a modest drop in classification accuracy ($2\%$) a ViT-Tiny model trained on the CIFAR-10 dataset when skip connections are removed from the FFN while retained in the SAB. However, removing skip connections from the SAB, while keeping them in the FFN, leads to a near-catastrophic performance drop of approximately $22\%$. Moreover, as dataset size increases, models without SAB skip connections degrade substantially in performance; on larger-scale datasets such as Tiny-ImageNet, these models fail to train altogether. To the best of our knowledge, this empirical finding has not been reported in prior literature. Additionally, the training loss curves for the three configurations are shown in Fig.˜2. The results demonstrate that the SAB without skip connections, which is significantly ill-conditioned, exhibits much slower convergence and diverges after 30 epochs. Interestingly, this critical dependence on skip connections is unique to transformer architectures.
@@
Skip in ConvMixer. To further highlight the poor conditioning of the SAB, we conducted a similar experiment on a modern CNN ConvMixer [28]. ConvMixer has a similar architecture and competitive performance to ViT with the exception that self-attention is replaced by a convolution block to spatially mix tokens. Unlike ViT, training ConvMixer-Tiny with and without skip connections, we observed that the performance remains effectively unchanged, with variations of less than ±0.2%. Further details are shown in Fig.˜6 in Appendix. This contrast further suggests that the SAB introduces an extreme conditioning issue, making skip connections indispensable for stable training in ViTs.
Our main contributions are summarized as follows.
@@
    We present a proposition that characterizes why SAB output embedding without skip connection is fundamentally ill-conditioned, which challenges training convergence and stability (Sec.˜4.1).
@@
    A theoretical analysis on the role of skip connection within the SAB is undertaken. We demonstrate that it significantly improves the condition of the block’s output embedding, enhancing stability and performance (Sec.˜4.2).
@@
    Finally, we propose a novel approach—Token Greying (TG)—to better pre-condition the input tokens. This step, when used in conjunction with conventional skip connections—improves ViT in both supervised and self-supervised settings (Sec.˜5).
@@ 2 Related work
Originally proposed by Vaswani et al[29], the self-attention mechanism has become central to transformer architectures, enabling the capture of long-range dependencies by dynamically weighting interactions between tokens because it captures long-range dependencies by dynamically weighting interactions between tokens. Building on this foundation, transformer-based large language models, such as Llama [6] and Deepseek [15], have significantly advanced language understanding. In computer vision, researchers introduced Vision Transformers (ViTs) [5] for image classification, achieving superior results compared to traditional convolutional neural networks. Moreover, researchers have trained Diffusion Transformers [21], replacing the commonly used U-Net backbone with a transformer that operates on latent patches and inherits the excellent scaling properties of the transformer model class. Similarly, self-attention mechanisms have been successfully applied in speech recognition and physics-informed neural networks, further demonstrating their versatility across diverse applications [3, 32].
@@
He et al[11] introduced skip connections in ResNet to make deep networks easier to optimize and to enhance accuracy by enabling significantly increased network depth. Since their introduction, skip connections have been extensively studied and applied across various Convolutional Neural Network (CNN) architectures. Building on these ideas, Veit et al[30] offer a novel interpretation of residual networks, demonstrating that they can be viewed as a collection of multiple paths of varying lengths, which contributes to the network’s flexibility and resilience. Furthermore, Hayou et al[9] addresses gradient stability in deep ResNets, noting that while skip connections alleviate the issue of vanishing gradients, they may also introduce gradient explosion at extreme depths. To counter this, they propose scaling the skip connections according to the layer index, thereby stabilizing gradients and ensuring network expressivity even in the limit of infinite depth. Recently, in Table 3 of [16], the authors reported that VGG—a model without skip connections—achieves comparable results to ResNet—a model with skip connections—on modern datasets. To the best of our knowledge, only one paper challenges the skip connection in transformer architectures. In [10], the authors attempted to train deep transformer networks without using skip connections, achieving performance comparable to standard models. However, this approach requires five times more iterations, and the reasons why skip connections are crucial for self-attention-based transformers remain unresolved.
@@ 3 Preliminary
In this section, we define the self-attention blocks used in the Vision Transformer (ViT) architecture and establish the mathematical notation for key quantities referenced in subsequent sections. For a comprehensive overview of transformers, we refer readers to [5].
A Vision Transformer architecture consists of L stacked self-attention blocks (SABs) and feedforward networks (FFNs). An identity mapping, commonly referred to as a skip connection, is applied to connect the inputs and outputs of the transformation in both SAB and FFN.
@@
where the self-attention output embedding is
@@
where $\mathbf{Q}=\mathbf{X}_{\text{in}}\mathbf{W}_{\text{Q}}$, $\mathbf{K}=\mathbf{X}_{\text{in}}\mathbf{W}_{\text{K}}$ and $\mathbf{V}=\mathbf{X}_{\text{in}}\mathbf{W}_{\text{V}}$ with learnable parameters $\mathbf{W}^{\text{Q}},\mathbf{W}^{\text{K}},\text{and }\mathbf{W}^{\text{V}}\in\mathbb{R}^{d\times d}$. $\eta(\cdot)$ is an activation function and is set by default as $\operatorname*{softmax}$ function in [5]. Additionally, some studies [26, 22, 23] have explored the possibilities of using alternative activation functions. In particular, good performance is achieved by using linear attention with an appropriate scale. It allows us to characterize why the SAB output embedding without a skip connection is ill-conditioned.
@@
where $\|\cdot\|_{2}$ is the matrix operator norm induced by the Euclidean norm. We use the Euclidean norm in this paper. $\mathbf{A^{\dagger}}$ is the pseudo inverse of the matrix $\mathbf{A}$. $\sigma_{\text{max}}$ and $\sigma_{\text{min}}$ are the maximal and minimal singular values of $\mathbf{A}$ respectively. We will use Eq.˜4 as a metric for analysis in future sections.
@@ 4.1 Self-attention is ill-conditioned
Assume $\mathbf{X}\in\mathbb{R}^{n\times d}$$\mathbf{W_{Q}}$, and $\mathbf{W_{K}}$ and $\mathbf{W_{V}}\in\mathbb{R}^{d\times d}$ have entries that are independently drawn from a distribution with zero mean. The condition number of the SAB output embedding (without skip-connection) can be expressed as:
@@
Fig.˜3 illustrates our theoretical analysis by showing that the linear self-attention operation increases the condition number of embeddings from $e^{3}$ to $e^{6}$, and we can empirically extend this analysis to softmax self-attention. The output embedding of the softmax SAB without skip connections is also significantly ill-conditioned.
@@ 4.2 Skip connection improves condition
Previous works [10, 18, 4] have demonstrated that transformers, without skip connections, converge very slowly or may even become un-trainable. In [10], the authors show that without skip connections, output embeddings of the SAB quickly converge to rank 1. They hypothesize that this is due to the inability of the transformer to propagate signals through the self-attention mechanism. This rank-1 output embedding implies an infinite condition number. The authors introduce an inductive bias into the self-attention mechanism to enable what they refer to as “better signal propagation”. While their approach matches the performance of standard (vanilla) models, it incurs a substantial computational cost—running approximately five times slower. We hypothesize that the success of this method could be directly attributed to its ability to improve the condition of the self-attention output embeddings.
This observation further motivates our view that the primary role of skip connections may be to regularize the condition of the self-attention output embeddings
@@
Proposition˜4.2 shows that during training, the condition number of the SAB output embedding has a much lower bound than that of the SAB output embedding without skip connections. Furthermore, with other conditions being equal, better condition of SAB can contribute to faster convergence and more stable gradient updates, reducing the risk of gradient explosion or vanishing. This ensures the signal can propagate well during the training process in the Vision Transformers.
@@
As shown in Fig.˜3, empirical results demonstrate that in both the linear and $\operatorname*{softmax}$ self-attention cases, adding skip connections significantly improves the condition of output embeddings.
@@ 5 Methodology
[…]
@@
5:  $\mathbf{U\Sigma V^{T}\leftarrow X}$ $\triangleright$ Compute SVD
6:  $\mathbf{\tilde{\Sigma}\leftarrow\frac{\Sigma}{\text{max}(\Sigma)}}$ $\triangleright$ Normalize elements $\in(0,1]$
7:  $\mathbf{\tilde{\Sigma}\leftarrow\tilde{\Sigma}^{\epsilon}}$ $\triangleright$ Amplify
8:  $\mathbf{\tilde{X}\leftarrow U\tilde{\Sigma}V^{T}}$ $\triangleright$ SVD reconstruct
[…]